A train is moving along a straight track with an initial velocity of 8 m/s when it is subjected to a deceleration of a-1.6/v m/s2, where v is in m/s. 2. How far does the train travel before it stops? How much time does it take to stop? a. b.
Answer #1

To calculate a.) we will use the formula

v2 - u2 = 2*a*s where v is final velocity, u is initial velocity and s is the distance.

For b.) we will use the formula

v = u +at

Given Sitia veloity u 8 )s final veloutyv deceleration a ) distance --t av el-by-train-一before (a -find t stopi ca -to s dist